OpenStudy (anonymous):

solve 8x - (3x+7) = 13

OpenStudy (anonymous):

8x-3x-7=13 5x=20 x=4

OpenStudy (lalaly):

5x=20 x=4

OpenStudy (anonymous):

In general, bring your unknowns (x's) to one side of the = and the knowns (integers) to the other side, which gives 8x - 3x = 13 + 7 or 5x = 20 and x = 4

OpenStudy (anonymous):

8x-3x-7=13 5x=20 x=4

OpenStudy (anonymous):

8x-3x-7=13 5x=13+7 5x=20 x=4